1. Find the general solution of the given higher-order differential equation.
16 d^4 y/ dx^4 + 24 d^2 y/ dx^2+ 9y = 0

y(x) =_____

2. Find the general solution of the given second-order differential equation.
y'' − 10y' + 26y = 0

y(x) = ______
1
Expert's answer
2020-06-28T16:41:46-0400

1) Given differential equation is "16 \frac{d^4 y}{dx^4} + 24 \frac{d^2 y}{dx^2}+ 9y = 0" .

"\implies (16D^4+24D^2+9)y = 0" where "D = \frac{d}{dx}" .

"\implies (4D^2+3)^2 y = 0"

So, Auxiliary equation is "(4D^2+3)^2=0 \implies D = \pm \frac{\sqrt{3}}{2}i, \pm \frac{\sqrt{3}}{2}i" .

Hence, "y(x) = ( c_1 +c_2 x) cos(\frac{\sqrt{3}}{2} x) + (c_3 +c_4 x) sin(\frac{\sqrt{3}}{2} x)" .


2) Given "y'' \u2212 10y' + 26y = 0"

"\implies (D^2-10D+26)y = 0" where "D = \frac{d}{dx}" .

So, Auxiliary equation is "D^2-10D+26=0"

"\implies D = \frac{10\pm \sqrt{100-104}}{2} = 5 \pm i" .

Hence, "y = e^{5x}(c_1cos(x)+c_2 sin(x))"